The perimeter of a window is 38 feet. The width of the window is three feet more than the length. What are the dimensions of the window?
Answer by cleomenius(959)

Perimeter = 2L + 2S = 38 feet.
2(x + 3) + 2x = 38
4x + 6 = 38.
4x = 32
x = 8 feet, the length.
x + 3 = 11 feet, the width.
22 + 16 = 38 feet.
Cleomenius.